Write the equation of the line using this information
a ) slope=2, y-intercept 0
b) slope 6, passing through (-4,7)
please help me with this equation
Thanks

Answer by bucky(2189)

Write the equation for a graph that has a slope = 2 and a y-intercept of 0
.....
The definition of the slope-intercept form of an equation is y = mx + b
in which
m is the slope and b is the y-intercept.
All you have to do now is substitute the given values of +2 for m and 0 for +b. This makes
the equation that you are looking for: y = +2x + 0 which reduces to just y = +2x
........
Write the equation for a graph that has a slope of 6 and passes through (-4,7).
You can again use the slope-intercept form:
y = mx + b
Substitute the given slope for m to get:
y = 6x + b
You also given that when x = -4 and y = +7 it satisfies the equation. Substituting
these two values into the equation result in:
7 = (6)*(-4) + b
After the multiplication you have:
7 = -24 + b
Solve this equation for b by adding 24 to both sides to get:
31 = 0 + b
which becomes just b = +31
Return to the equation y = 6x + b and substitute +31 for b. When you do the equation
that has a slope of 6 and goes through the point (-4,+7) becomes:
y = 6x + 31
Hope this gives you a little insight into graphing equations of the slope-intercept
form.